Baba Rafi is considering opening a small sandwich outlet inside the NSU campus. It will require  an initial investment of $20,000 and throughout the next 5 years the project will potentially  generate free cash flows in the following form: 

5

-20,000 

6000 

10000 

-4000 

3500 

6500



Now, as we can see an unconventional cash flow in year 3, please calculate the MIRR to decide  whether Baba Rafi should invest in this project or not? The required rate of return is the WACC  calculated in problem 4.




1
Expert's answer
2022-02-03T14:34:21-0500


let WACC 5%

First, we will discount the positive cash flows

"\frac{6000}{(1+0.05)^1}+\frac{10000}{(1+0.05)^2}+\frac{3500}{(1+0.05)^4}+\frac{6500}{(1+0.05)^5}=22756.96"


Then we discount the negative cash flow

"20000-\frac{-4000}{(1+0.05)^3)}=16544.65"

"MIRR=(\frac{22756.96}{16544.65})^\frac{1}{5}-1=0.07" or 7%


the project is the project is recommended for implementation MIRR >WACC